# Qudit inspired optimization for graph coloring

### v1

@@ III Benchmark
Figure 3: (a) Average of color conflicts corresponding to the Potts energy from Eq. (1) for the queen11-11 graph as a function of total number of time steps $N_{steps}$. We use $c=11$ colors, $f=0.0$, and different $h$. For the calculations we use $N_{\rm{runs}}=100$. The black dashed line is at 13, the best solution presented in Ref. wang_23 . (b) Standard deviation for the data shown in (a). (c) Histogram showing the number, $N_{\textrm{conf}}$, a certain energy was found for $N_{steps}=1000$. Also here, the dashed line is at 13. All parameters not specified are provided in Appendix A.
Figure 4: (a) Total energy and Potts energy for different values of the learning rate $\eta$ for $N_{steps}=1000$ and $c=11$. The minimum values of $E_{\textrm{Potts}}$ reached in the simulations were 19, 11 (10), and 12 for $\eta=0.01$, 0.5 ($f=0.1$), and 1.0 respectively (we use $N_{\textrm{runs}}=100$). (b) Average number of color clashes [$E_{\textrm{Potts}}$ from Eq. (1)] for the same parameters as (a). The orange line additionally shows the data calculated using QdGD. All other parameters can be found in Appendix A.
We benchmark our algorithms by considering several standard GCP instances used to test other methods hebrard_19 ; li_22 ; schuetz_22_2 ; wang_23 ; brighen_23 and comparing to results in the literature hebrard_19 ; schuetz_22_2 ; wang_23 . First, we consider Myciel and Queen graphs from the COLOR data set colordataset_02 and the citation networks Cora and Pubmed McCallum_2000 ; Sen_2008 ; nr-aaai15 . The Myciel and Queen graphs, based on Mycielski transformation and $n\times m$ chessboards respectively, have up to 169 nodes and densities between $16.91$ and $53.33\%$. In contrast, the Cora and Pubmed graphs have 2708 and 19717 nodes and densities of $0.15\%$ and $0.02\%$. Note that a more thorough description of the datasets can be found in Ref. schuetz_22_2 .
We then consider three graphs from the SNAP library snapnets : The Facebook dataset based on social circles from Facebook, the Wikipedia vote dataset based on Wikipedia voting data, and the email-Eu-core data set based on Email communication between members of a research institute. For these three graphs, it is known hebrard_19 that relatively large numbers of colors are required to solve the problems (we use 70, 22, and 19) and thus they provide interesting challenges for our algorithms. In all cases, we consider undirected graphs, and we remove duplicated edges and any non-connected nodes when prepossessing the data.
@@ III.1 Influence of the hyperparameters
First, we illustrate how some of the hyperparameters in the QdLQA algorithm influence the final solutions. As an example, we show results for the queen11-11 graph. In Fig. 3(a), we demonstrate how the average value of the Potts energy depends on the number of time steps, $N_{\textrm{steps}}$, for different values of $h$ [which generates randomness in $E_{\textrm{F}}(\theta)$ in Eq. (9)]. Increasing the number of steps improves the average quality of the solutions. The black dashed line shows the best value for $E_{\rm{Potts}}$ reported in Ref wang_23 . Similar behaviour is seen for the standard deviation of the $h=3$ data in Fig. 3(b). Additionally, we see that for large $N_{\textrm{steps}}$, the $h=10$ data has a larger standard deviation and average value than $h=3$. We observe that having a finite $h$ helps the algorithm escape from local minima, but if $h$ is chosen too large, this comes at the cost of more low-quality solutions. This is further illustrated in Fig. 3(c), where we show a histogram of the number of times, $N_{\rm{conf}}$, a configuration with a certain energy is found. Again, the black dashed line shows the best value for $E_{\rm{Potts}}$ reported in Ref wang_23 . There, we see how also choosing $h=2$ increases the fluctuations and leads to the algorithm getting stuck more frequently in a bad local minima. In 42 instances, the solution found by $h=2$ was out of the scale of the figure. Correspondingly, the $h=2$ data are also not plotted in Fig. 3(a).
@@
We now show an example of how the cost function $E_{\rm{Total}}$ and the Potts energy $E_{\rm{Potts}}$ behave during the quantum adiabatic evolution in QdLQA. In Fig. 4, we plot both for the queen11-11 graph as a function of $t$ averaged over $N_{\textrm{runs}}=100$ runs for different learning rates $\eta$. For $E_{\rm{Total}}$ [Fig. 4(a)], we see that a large learning rate leads to an initial jump in the cost function, which increases when $\eta$ is increased. We attribute this jump to the fact that we start in the global minima at $t=0$, which remains close to the minima for small increments in $t$. Thus, doing a gradient descent step with a large learning rate will make us "overshoot" and take us away from the presumptive minima. However, we see that the $\eta=0.5$ curve quickly relaxes to the $\eta=0.01$, whereas the $\eta=1.0$ curve also shows large fluctuations for later times. Note that in contrast to most schemes using the Adam optimizer, here, it depends on the energy landscape at previous times and the varying cost function [due to the randomness introduced by $h$ in Eq. (9)]. By setting $f=0.1$ (black dashed line), we observe that the amplitude of the initial jump increases, but that the curve soon follows the $f=0.0$ curve.
@@
Figure 4(b) shows $E_{\rm{Potts}}$ for the same parameters as Fig. 4(a). Here, one can see large fluctuations in the energies for $\eta=0.5$ and $\eta=1.0$, but which decrease when $t$ is increased. The initial decay coincide with the jump in Fig. 4(a) and can be understood as follows: In the initial state, all qudits are initialized equally, and thus, $E_{\rm{Potts}}$ has maximum value. If one now goes away from the minimum, this leads to different qudits taking on different colors, and thus the value of the final cost function will decrease. This can also be seen by changing the perturbation of the initial state (setting $f=0.1$ is illustrated by the black dashed line). Now, the qudit states on different nodes differ, and thus $E_{\rm{Potts}}$ is smaller at $t=0$. The orange curve in Fig. 4(b) shows $E_{\rm{Potts}}$ for the QdGD algorithm (as a function of gradient descent steps $n=N_{\textrm{steps}}t$). Notably, this decays rapidly, which allows for a good solution to be found much earlier. When stopping the QdGD after having 100 gradient descent steps without improving $E_{\rm{Potts}}$, the average number of steps was $n=259.28$ (corresponding to $t\approx 0.26$).
@@ III.2 Comparing with other methods
Table 1: Datasets from colordataset_02 and the citation graphs McCallum_2000 ; Sen_2008 ; nr-aaai15 introduced in the main text. Here, “colors” represents the known minimum number of colors required to color the graph, whilst the following values are related to the cost function obtained after optimization by various methods, such that zero is optimum. Values given for Tabucol schuetz_22_2 , PI-CGN schuetz_22_2 , PI-SAGE schuetz_22_2 and GNN-1N wang_23 represent other algorithms, whilst the quantities related to $E^{\textrm{QdLQA}}_{\text{Potts}}$ and $E^{\textrm{QdGD}}_{\text{Potts}}$ show the best solution found by the corresponding method in $N_{\textrm{runs}}$ runs. The parameters $p^{\textrm{QdLQA}}_{\textrm{min}}$ and $p^{\textrm{QdLQA}}_{\textrm{min}}$ give the fractions, $N_{\textrm{min}}/N_{\textrm{runs}}$, indicating how often, $N_{\textrm{min}}$, the solution was found in the $N_{\textrm{runs}}$ runs. The best solution found by any of the algorithms is indicated in bold font, NA means that no solution was found within 24 hours schuetz_22_2 , and X means no value was provided in Ref. wang_23 . For Cora and Pubmed, the choice of $c$ follows Refs. li_22 ; schuetz_22_2 .
We now compare the solutions found by our algorithms to previously published results using graph neural networks in Refs. schuetz_22_2 ; wang_23 . In Table 1, we show the energies of the best configurations found with QdLQA and QdGD ($E^{QdLQA}_{\rm{Potts}}$ and $E^{QdGD}_{\rm{Potts}}$ respectively) and the fraction of the runs that found that energy ($p^{QdLQA}_{\textrm{min}}$ and $p^{QdGD}_{\textrm{min}}$) for different graphs. Furthermore, we show the corresponding values reported in Refs. schuetz_22_2 ; wang_23 obtained using a Tabucol hertz_87 and a greedy algorithm, a physics-inspired graph neural network solver (PI-SAGE) schuetz_22_2 , and graph neural networks using a first-order negative message passing strategy (GNN-1N) wang_23 . Note that if $E_{\rm{Potts}}=0$, the graph was colored with no conflicts, else, $E_{\rm{Potts}}>0$ gives the remaining number of color clashes. For all the graphs except queen11-11, queen13-13, Cora, and PubMed, both our algorithms find the optimal solution. however, some graphs seem to be particularly hard. For example, for the queen8-8 and queen9-9 graphs, both algorithms only find the optimal solutions in $<10$ of the 100 runs. For queen11-11 and queen13-13, QdLQA finds better solutions than those reported in Refs. schuetz_22_2 ; wang_23 . In the case of queen11-11, it additionally finds $E_{\rm{Potts}}=11$ in three of the 100 runs, and both $E_{\rm{Potts}}=12$ and $E_{\rm{Potts}}=13$ in nine of the 100 runs each. For queen13-13, it additionally finds $E_{\rm{Potts}}=13$ in two, $E_{\rm{Potts}}=14$ in two, and $E_{\rm{Potts}}=15$ in seven of the 100 runs.
@@
For the Cora graph, only the QdGD algorithm finds the $E_{\rm{Potts}}=0$ solution, and only in one of the 100 runs. QdLQA finds $E_{\rm{Potts}}=1$ configurations in three and $E_{\rm{Potts}}=2$ in 20 of the 100 runs, but requires a comparably large $h$ to do so (we use $h=10$). In this case, one must compromise between finding $E_{\rm{Potts}}=1$ a few times at the cost of obtaining many bad solutions. For example, using $h=3$ instead, we get $E_{\rm{Potts}}=2$ in 58 of 100 runs, but never $E_{\rm{Potts}}=1$. We also note that, as shown in Appendix A, the average runtime for one QdGD run on a laptop is significantly less than what is reported for PI-GCN in Ref. schuetz_22_2 for the same graphs on p3.2×large AWS EC2 instances.
For the PubMed graph, neither algorithm finds better solutions than the best one reported in Ref. schuetz_22_2 , but QdLQA seems to do significantly better than QdGD. Additionally, in contrast to the other datasets in Table 1, we do not fix the node with the largest number of edges for the PubMed graph. Rather, we fix a node with one edge as this improves the results. For $N_{\textrm{runs}}=100$, QdLQA improves from finding $E^{QdLQA}_{\textrm{Potts}}=17$ three times to finding $E^{QdLQA}_{\textrm{Potts}}=15,$ $16,$ and $17$ three, one, and seven times respectively. QdGD improves from finding $E^{QdGD}_{\textrm{Potts}}=28$ twice to finding $E^{QdGD}_{\textrm{Potts}}=27$ three times. This indicates that cleverly fixing one (or more) nodes might lead to further improvements.
Note that we provide a comprehensive overview of the average runtimes and computer architectures used in Appendix A and a list of parameters used in Appendix B.
@@ III.3 Benchmarking using large number of colors
Finally, we test our algorithms on graphs where it is known that relatively large numbers of colors are needed to solve the problems hebrard_19 . In Table 2, we show results for the previously mentioned email-Eu-core, Facebook, and Wiki-vote datasets together with the corresponding upper and lower bounds provided in Ref. hebrard_19 (the problem was solved for email-Eu-core and Facebook). In our simulations, we use the number of colors equal to the upper bounds provided in Ref. brighen_23 (19 for email-Eu-core, 70 for Facebook, and 22 for Wiki-vote), and again, $E^{QdLQA}_{\rm{Potts}}$ and $E^{QdGS}_{\rm{Potts}}$ give the number of conflicts in the best solution found. Even though neither algorithm solves the problem completely, they are clearly capable of providing relatively good solutions (quantified by the normalized error defined as the energy divided by number of edges, $\epsilon=E_{\rm{Potts}}/\absolutevalue{\mathcal{E}}$ | schuetz_22_2 ). For the email-Eu-core, Facebook, and Wiki-vote datasets, $\epsilon\approx 0.0016$ (0.0019), 0.00028 (0.00025) and 0.0019 (0.002) for QdLQA (QdGS). This demonstrates the ability of the algorithms to find good solutions for relatively large graphs requiring many colors on a simple laptop or desktop (see Appendix A for details).

### v2

@@ III Benchmark
We benchmark our algorithms by considering several standard GCP instances used to test other methods hebrard_19 ; li_22 ; schuetz_22_2 ; wang_23 and comparing to results in the literature hebrard_19 ; schuetz_22_2 ; wang_23 . First, we consider Myciel and Queen graphs from the COLOR data set colordataset_02 and the citation networks Cora and PubMed McCallum_2000 ; Sen_2008 ; nr-aaai15 . The Myciel and Queen graphs, based on Mycielski transformation and $n\times m$ chessboards respectively, have up to 169 nodes and densities between $16.91\%$ and $53.33\%$. In contrast, the Cora and PubMed graphs have 2708 and 19717 nodes and densities of $0.15\%$ and $0.02\%$. Note that a more thorough description of the datasets can be found in Ref. schuetz_22_2 .
We then consider three graphs from the SNAP library snapnets : The Facebook dataset based on social circles from Facebook, the Wikipedia vote dataset based on Wikipedia voting data, and the email-Eu-core data set based on email communication between members of a research institute. For these three graphs, it is known hebrard_19 that relatively large numbers of colors are required to solve the problems (we use 70, 22, and 19) and thus they provide interesting challenges for our algorithms. In all cases, we consider undirected graphs, and we remove duplicated edges and any nonconnected nodes when prepossessing the data.
@@ III.1 Influence of the hyperparameters
Figure 3: (a) Average of color conflicts corresponding to the Potts energy from Eq. (1) for the queen11-11 graph as a function of total number of time steps $N_{\text{steps}}$. We use $c=11$ colors, $f=0.0$, and different $h$. For the calculations we use $N_{\rm{runs}}=100$. The black dashed line is at 13, the best solution presented in Ref. wang_23 . (b) Standard deviation for the data shown in (a). (c) Histogram showing the number, $N_{\textrm{conf}}$, of times a certain energy was found for $N_{\text{steps}}=1000$. Also here, the dashed line is at 13. All parameters not specified are provided in Appendix C.
Figure 4: (a) Total energy for different values of the learning rate $\eta$ for $N_{\text{steps}}=1000$ and $c=11$. (b) Same data as in (a) but on a shorter time scale and showing the individual data points. All other parameters can be found in Appendix C.
Figure 5: Average number of color clashes [$E_{\textrm{Potts}}$ from Eq. (1)] for the same parameters as in Fig. 4. The orange line additionally shows the data calculated using QdGD and the minimum values of $E_{\textrm{Potts}}$ reached in the simulations were 19 ($\eta=0.01$, $f=0.0$), 11 ($\eta=0.5$, $f=0.0$), 10 ($\eta=0.5$, $f=0.1$), 12 ($\eta=0.01$, $f=0.0$), and 13 (QdGD) using $N_{\textrm{runs}}=100$. All other parameters can be found in Appendix C.
@@
First, we illustrate how some of the hyperparameters in the QdLQA algorithm influence the final solutions. As an example, we show results for the queen11-11 graph. In Fig. 3(a), we demonstrate how the average value of the Potts energy depends on the number of time steps, $N_{\textrm{steps}}$, for different values of $h$ [which generates randomness in $E_{\textrm{F}}({\psi})$ in Eq. (9)]. Increasing the number of steps improves the average quality of the solutions. The black dashed line shows the best value for $E_{\rm{Potts}}$ reported in Ref wang_23 . Similar behavior is seen for the standard deviation of the $h=3$ data in Fig. 3(b). In addition, we see that for large $N_{\textrm{steps}}$, the $h=10$ data have a larger standard deviation and average value than $h=3$. We observe that having a finite $h$ helps the algorithm escape from local minima, but if $h$ is chosen too large, this comes at the cost of more low-quality solutions. This is further illustrated in Fig. 3(c), where we show a histogram of the number of times, $N_{\rm{conf}}$, a configuration with a certain energy is found. Again, the black dashed line shows the best value for $E_{\rm{Potts}}$ reported in Ref wang_23 . There, we see how also choosing $h=2$ increases the fluctuations and leads to the algorithm getting stuck more frequently in a bad local minimum. In 42 instances, the solution found by $h=2$ was outside the scale of the figure. Thus, the $h=2$ data are also not plotted in Fig. 3(a).
@@
We now show an example of how the cost function $E_{\rm{Total}}$ and the Potts energy $E_{\rm{Potts}}$ behave during the quantum adiabatic evolution in QdLQA. In Fig. 4(a), we plot $E_{\rm{Total}}$ for the queen11-11 graph as a function of $t$ averaged over $N_{\textrm{runs}}=100$ runs for different learning rates $\eta$. There, we see that a large learning rate leads to an initial jump in the cost function, which increases when $\eta$ is increased. This can be seen better in Fig. 4(b), where the individual data points of the initial dynamics are plotted. We attribute this jump to the fact that we start in the global minimum at $t=0$, which remains close to the minimum for small increments in $t$. Thus, doing a gradient descent step with a large learning rate will make us "overshoot" and take us away from the presumptive minima. However, we see that the $\eta=0.5$ curve quickly relaxes to the $\eta=0.01$ curve for a large part of the time evolution. In contrast, the $\eta=1.0$ curve also shows large fluctuations for later times.
@@
Figure 5 shows $E_{\rm{Potts}}$ for the same parameters as Fig. 4. Here, one can see large fluctuations in the energies for $\eta=0.5$ and $\eta=1.0$, but which decrease when $t$ is increased. The initial decay coincides with the jump in Fig. 4 and can be understood as follows: In the initial state, all qudits are initialized equally, and thus, $E_{\rm{Potts}}$ has maximum value. If one now goes away from the minimum, this leads to different qudits taking on different colors, and thus the value of $E_{\rm{Potts}}$ will decrease. If one perturbs the initial state (the $f=0.1$ setting is illustrated by the black dots), the colors at different qudits will differ, and thus, $E_{\rm{Potts}}$ is smaller at $t=0$. However, at short times, the cost function will still favor all colors being the same [the ground state of Eq. (8)], and thus $E_{\rm{Potts}}$ increases. For larger times, it then closely resembles the curve for the unperturbed initial state (blue solid line) as it decreases. The orange curve in Fig. 5 shows $E_{\rm{Potts}}$ for the QdGD algorithm (as a function of gradient descent steps $n=N_{\textrm{steps}}t$). Notably, this decays rapidly, which allows for a good solution to be found much earlier. When stopping the QdGD after having 100 gradient descent steps without improving $E_{\rm{Potts}}$, the average number of steps was $n=259.28$ (corresponding to $t\approx 0.26$).
Lastly, we want to point out another interesting feature of our approach. In most applications (e.g., often when training neural networks), optimizers like Adam update their internal parameters based on previous gradient descent steps for a fixed cost function. Here, in contrast, the parameters are updated based on the energy landscape at previous times (due to the varying cost function).
@@ III.2 Comparing with other methods
Table 1: Datasets from colordataset_02 and the citation graphs McCallum_2000 ; Sen_2008 ; nr-aaai15 introduced in the main text. Here, “Colors” represents the number of colors used in the calculations, while the following values are related to the cost function obtained after optimization by various methods, such that zero is optimum. Values given for Tabucol schuetz_22_2 , PI-CGN schuetz_22_2 , PI-SAGE schuetz_22_2 , and GNN-1N wang_23 represent other algorithms, whilst the quantities related to $E^{\textrm{QdLQA}}_{\text{Potts}}$ and $E^{\textrm{QdGD}}_{\text{Potts}}$ show the best solution found by the corresponding method in $N_{\textrm{runs}}$ runs. The parameters $p^{\textrm{QdLQA}}_{\textrm{min}}$ and $p^{\textrm{QdLQA}}_{\textrm{min}}$ give the fractions, $N_{\textrm{min}}/N_{\textrm{runs}}$, indicating how often, $N_{\textrm{min}}$, the solution was found in the $N_{\textrm{runs}}$ runs. The best solution found by any of the algorithms is indicated in bold face, NA means that no solution was found within 24 h schuetz_22_2 , and X means no value was provided in Ref. wang_23 . For Cora and PubMed, the choice of $c$ follows Refs. li_22 ; schuetz_22_2 . * 17 was reported in schuetz_22_2 but we find 13 in our own calculations, see Appendix B.
We now compare the solutions found by our algorithms to previously published results using graph neural networks in Refs. schuetz_22_2 ; wang_23 . In Table 1, we show the energies of the best configurations found with QdLQA and QdGD ($E^{QdLQA}_{\rm{Potts}}$ and $E^{QdGD}_{\rm{Potts}}$, respectively) and the fraction of the runs that found that energy ($p^{QdLQA}_{\textrm{min}}$ and $p^{QdGD}_{\textrm{min}}$) for different graphs. Furthermore, we show the corresponding values reported in Refs. schuetz_22_2 ; wang_23 obtained using a Tabucol hertz_87 and a greedy algorithm, a physics-inspired graph neural network solver (PI-SAGE) schuetz_22_2 , and graph neural networks using a first-order negative message passing strategy (GNN-1N) wang_23 . Note that if $E_{\rm{Potts}}=0$, the graph was colored with no conflicts, otherwise, $E_{\rm{Potts}}>0$ gives the remaining number of color clashes. For all the graphs except queen11-11, queen13-13, Cora, and PubMed, both our algorithms find the optimal solution. However, some graphs seem to be particularly hard. For example, for the queen8-8 and queen9-9 graphs, both algorithms only find the optimal solutions in less than $10$ of the 100 runs. For queen11-11 and queen13-13, QdLQA finds better solutions than those reported in Refs. schuetz_22_2 ; wang_23 . In the case of queen11-11, it additionally finds $E_{\rm{Potts}}=11$ in three of the 100 runs, and both $E_{\rm{Potts}}=12$ and $E_{\rm{Potts}}=13$ in nine of the 100 runs each. For queen13-13, it additionally finds $E_{\rm{Potts}}=13$ in two, $E_{\rm{Potts}}=14$ in two, and $E_{\rm{Potts}}=15$ in seven of the 100 runs.
@@
For the Cora graph, only the QdGD algorithm finds the $E_{\rm{Potts}}=0$ solution, and only in one of the 100 runs. QdLQA finds $E_{\rm{Potts}}=1$ configurations in three and $E_{\rm{Potts}}=2$ in 20 of the 100 runs, but requires a comparatively large $h$ to do so (we use $h=10$). In this case, one must compromise between finding $E_{\rm{Potts}}=1$ a few times at the cost of obtaining many bad solutions. For example, using $h=3$ instead, we get $E_{\rm{Potts}}=2$ in 58 of 100 runs, but never $E_{\rm{Potts}}=1$.
For the PubMed graph, neither algorithm finds better solutions than the best one reported in Ref. schuetz_22_2 , but QdLQA seems to do significantly better than QdGD. In addtion, in contrast to the other datasets in Table 1, we do not fix the node with the largest number of edges for the PubMed graph. Rather, we fix a node with one edge as this improves the results. For $N_{\textrm{runs}}=100$, QdLQA improves from finding $E^{QdLQA}_{\textrm{Potts}}=17$ three times to finding $E^{QdLQA}_{\textrm{Potts}}=15,$ $16,$ and $17$ three times, onece, and seven times, respectively. QdGD improves from finding $E^{QdGD}_{\textrm{Potts}}=28$ twice to finding $E^{QdGD}_{\textrm{Potts}}=27$ three times. This indicates that cleverly fixing one (or more) nodes might lead to further improvements.
Note that we provide a comprehensive overview of the average runtimes, computer architectures used, and a more thorough comparison to PI-SAGE in Appendix B and a list of parameters used in Appendix C.
@@ III.3 Benchmarking using large number of colors
Finally, we test our algorithms on graphs where it is known that relatively large numbers of colors are needed to solve the problems hebrard_19 . In Table 2, we show results for the previously mentioned email-Eu-core, Facebook, and Wiki-vote datasets together with the corresponding upper and lower bounds provided in Ref. hebrard_19 (the problem was solved for email-Eu-core and Facebook). In our simulations, we use the number of colors equal to the upper bounds provided in Ref. hebrard_19 (19 for email-Eu-core, 70 for Facebook, and 22 for Wiki-vote), and again, $E^{QdLQA}_{\rm{Potts}}$ and $E^{QdGS}_{\rm{Potts}}$ give the number of conflicts in the best solution found. Even though neither algorithm solves the problem completely, they are clearly capable of providing relatively good solutions (quantified by the normalized error defined as the energy divided by number of edges, $\epsilon=E_{\rm{Potts}}/\absolutevalue{\mathcal{E}}$ | schuetz_22_2 ). For the email-Eu-core, Facebook, and Wiki-vote datasets, $\epsilon\approx 0.0016$ (0.0019), 0.00028 (0.00025) and 0.0019 (0.002) for QdLQA (QdGS). This demonstrates the ability of the algorithms to find good solutions for relatively large graphs requiring many colors on a simple laptop or desktop (see Appendix B for details).